HUBBLE SPACE TELESCOPE DAILY REPORT #1059 PERIOD COVERED: 0700 EST Daily Status Report as of 024/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 4892 (Duplicity Among Hyades Stars, Cycle 3 High) The FGSs were used to make astrometric observations of the possible binary star VA119 which is located in the Hyades open cluster. The OSS monitored the observation, and no anomalies were reported. 1.2 Completed Seven Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Field Calibration) The WF/PC-2 was used to make a set of visual internal flat images. 23 images were received and displayed. Seven observations were reported to be underexposed, and the other were okay. 1.3 Completed COSTAR/U 4745 (COSTAR/FOS Fine Alignment - Focus) The Faint Object Spectrograph was scheduled to be used to make alignment test of the COSTAR FOS M1 mirror. These observations were not needed, and were canceled. 1.4 Completed S/C 4616 (WF/PC-2 Fine Alignment and focus) These observations were not needed, and were canceled. 1.5 Completed Three Sets of S/C 5254 (GHRS SMOV Side One Activation) The Side One DEB was successfully turned on, at three hours intervals. No anomalies were noted. 1.6 Completed Two Sets of FOS/RD 4908 (SMOV FOS/FGS Fine Alignment) The Faint Object Spectrograph was used to find the location of its apertures, the plate scale, and the orientation and separation of its detectors. Each part was executed on both the Red and Blues Sides. A corrective SIP maneuver was successfully uplinked to the spacecraft. The subsequent observations were reported to look fine. 1.7 Completed Two Sets of WF/PC-2 4819 (WF/PC-2 PSF Characterization) The WF/PC-2 was used to make multiple exposures of the globular Cluster Omega Centauri during coarse and fine dwell scans. The fine scans were only 0.025 arcseconds apart. Two different filters were used. Two images were received and displayed. The images were reported to look fine, except that they were exposures of the wrong target. 1.8 Completed GHRS/2 4891 (GHRS/FGS Coarse Alignment) The High Resolution Spectrograph was used to make observations of the Small Magellanic Cloud cluster NGC-346. This is to help confirm the location of its aperture locations. Four observations were received and displayed. They were reported to look good. 1.9 Completed WF/PC-2 4721 (WF/PC-2 Decontamination and TEC Turn On) The WF/PC-2 was used to make ultraviolet photometric observations of the star standard A+81D266, plus K-Spots, and dark calibration images. All four sets of CCDs were used. A total of 11 images were received and displayed in the OSS. No anomalies were reported. 1.10 Completed FGS 5057 (Extragalactic Astrometry and Astrophysics, FT0-11, Part 1 of 2) Fine Guidance Sensor #3 was used to make astrometric observations of of object 0826-373IN and 1521-REF. These observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. This proposal is tied to proposal 1013. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed FGS 5063 (Linking the HIPPARCOS to the Extragalactic Reference Frame, Part 5 of 6) The Fine Guidance Sensors were used to make positional observations of the HIPPARCOS star NEWEGOC-45 relative to that of a faint extragalactic object (ERO) in order to connect the galactic reference system to the extragalactic system. The observations were monitored while telemetry was available, and no problems were noted. 1.12 Completed Two Sets of WF/PC-2 4749 (WF/PC-2 Relative Photometric Calibration) The WF/PC-2 was used to make photometric calibrations images of standard star A+81D266 which is located in NGC-5136 (aka, Omega Centauri) to obtain various medium and broad band filter images. The OSS monitored the observations and assessed the data quality. No anomalies were noted. 1.13 Completed WF/PC-2 4709 (WF/PC-2 Internal Monitor) The Faint Object Spectrograph (Blue and Red Sides) was used to make internal dark images to help in the instrument's calibration. The observations were received and displayed. The images were reported to be normal. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 18 Scheduled Reacquisitions : 17 Successful : 17 2.2 FHST Updates: Scheduled :32 Successful :30 Failed the second part of a pair Full Maneuvers at 022/2017Z. Failed a roll delay at 023/0004Z. 2.3 Significant HSTARs: None 2.4 Operations Notes: SMOV activities continue on or ahead of the milestone schedule. During the past week WF/PC-2 proceeded with its decontamination monitoring, performed flat fields calibration, and executed QE and and throughput measurements. FOC, operating with COSTAR, performed point spread function measurements, plate scale calibrations, and further fine alignment observations (using a different target star having better characteristics for aligning the FOS). Finally, the GHRS, utilizing the COSTAR, successfully deployed its M2 mirror arm. All HST subsystems, including those with replacement components, are performing nominally 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Four Sets of WF/PC-2 5240 (SMOV-WF/PC-2 Flat Fields Calibration) 3.2 Complete GHRS/2 4731 (COSTAR/GHRS Coarse Alignment) 3.3 Complete WF/PC-2 5590 (Cycle 4 DD HST Observations of Comet Shoemaker-Levy (1993E)) 3.4 Complete WF/PC-3 5475 (A Snapshot Survey for Blue Stragglers in Magellanic Cluster Cores) 3.5 Complete WF/PC-2 5248 (WF/PC-2 Image Quality and Stability Check) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-